# Feasibility of Learning Weighted Automata on a Semiring

### v3

@@ 2.2. Learning with membership and equivalence queries
Suppose that $Q\subseteq S^{\Sigma^{*}}$ is a finite set that is left-closed and left-generates $\langle\varepsilon\rangle$. By definition, $\Lambda_{Q}$ is non-empty. We prove that if $\lambda$ belongs to $\Lambda_{Q}$ then $\mathcal{H}_{Q,\lambda}$ computes $f$. We first show that for all $q$ in $Q$ and $w$ in $\Sigma^{*}$, $\langle q\rangle_{w}$ is equal to the “weight of $w$ from $q$” that is, the weight computed on input $w$ by the automaton obtained from $\mathcal{H}_{Q,\lambda}$ by imposing the condition that $q$ is the unique initial state with weight $1_{S}$, and keeping all other weights the same. The proof is by induction on the length of $w$. If $w=\varepsilon$, $\langle q\rangle_{\varepsilon}$ is (by construction) equal to the final weight of $q$ in $\mathcal{H}_{Q,\lambda}$ which proves the property. For $a$ in $\Sigma$ and $w$ in $\Sigma^{*}$, $\langle q\rangle_{aw}=\langle qa\rangle_{w}=\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes\langle p\rangle_{w}$ since $\lambda\in\Lambda_{Q}$. By induction hypothesis, $\langle p\rangle_{w}$ is the weight of $w$ from $p$, and moreover $\lambda_{q,a,p}$ is (by construction) the weight of the transition from $q$ to $p$ labelled by $a$. So $\langle q\rangle_{aw}$ is indeed the weight of $aw$ from $q$. To conclude the proof, recall that $\langle\varepsilon\rangle$ denotes the row of $\varepsilon$ in the Hankel matrix of $f$, and so by definition of the Hankel matrix of $f$ and the fact that $\lambda\in\Lambda_{Q}$ we have $f(w)=\langle\varepsilon\rangle_{w}=\bigoplus_{q\in Q}\lambda^{\varepsilon}_{q}\otimes\langle q\rangle_{w}$; by our previous observation the latter is then exactly the value computed by $\mathcal{H}_{Q,\lambda}$ on input $w$.
@@
Now suppose that $f:\Sigma^{*}\to S$ is computed by a finite-state weighted automaton $\mathcal{A}$ over $S$ with set of states $Q$. For each $q\in Q$, let ${}_{q}\mathcal{A}$ (resp. $\mathcal{A}_{q}$) denote the automaton obtained from $\mathcal{A}$ by making $q$ the unique initial (resp. final) state with weight $1_{S}$. Define $\langle q\rangle\in S^{\Sigma^{*}}$ where $\langle q\rangle_{u}$ is equal to the value computed by ${}_{q}\mathcal{A}$ on $u$. Let $R=\{\langle q\rangle\mid q\in Q\}$. First $R$ is left-closed since, by definition, for all words $u\in\Sigma^{*}$, all states $q\in Q$ and all $a\in\Sigma$ we have
$\langle qa\rangle_{u}=\langle q\rangle_{au}=\bigoplus_{p\in Q}(\nu_{p}\otimes\langle p\rangle_{u})$
@@
where $\nu_{p}$ is the weight of the transition labelled by $a$ in $\mathcal{A}$ from $q$ to $p$. Let $w$ be a word in $\Sigma^{*}$ and let $\langle w\rangle$ denote the infinite row of the Hankel matrix of $f$ indexed by $w$. Let $\mu_{q}(w)$ be the value computed by $\mathcal{A}_{q}$ on input $w$. Then, by definition, $\langle w\rangle=\bigoplus_{q\in Q}(\mu_{q}(w)\otimes\langle q\rangle)$, showing that $R$ is row-generating for $f$ and hence in particular $\langle\varepsilon\rangle$ is left-generated by $R$. ∎
@@
In this section we have taken an approach based on “rows”, describing how to construct hypothesis automata utilising finitely many partial rows of the Hankel matrix. Clearly, a dual construction can be made working with the transpose of the Hankel matrix, that is, using partial columns an corresponding notions of right-linear combinations, right semimodules, right-generating sets, right-closed sets, column-generating sets, etc. In the text that follows we shall refer to the analogues of Hankel automata and hypothesis automata constructed by this dual method as co-Hankel automata and co-hypothesis automata. For the readers convenience we outline the details of these constructions and summarise the dual definition to those presented in the main text in Appendix A.
@@
It is known that for all functions computed by weighted automata over fields there is an efficient way to find $Q$, $T$ and $\lambda$, by membership and equivalence queries such that the hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ computes the target function (see for example [8]). For general semirings, the following questions arise:
@@
  • For which functions $f:\Sigma^{*}\rightarrow S$ does there exist a hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ computing $f$? We will call the class of all such functions weakly guessable. Clearly, if $f$ is not weakly guessable then any learning algorithm based around making hypotheses of this kind will fail to learn $f$.
@@
  • For the class of weakly guessable functions, is there a learning algorithm using membership and equivalence queries that allows a learner to construct a finite sequence of hypothesis automata, where the last in the sequence computes the target function? In other words: when there exists at least one correct hypothesis, can we devise a strategy to find one? A possible issue here is that even when we have settled upon a suitable set $Q$ with $\Lambda_{Q}\neq\emptyset$, it may be that there is no finite set $T$ such that $\Lambda_{Q,T}\setminus\Lambda_{Q}$ is finite, and there is no obvious general method to discover a “correct” $\lambda$ from the coefficient set in finite time. We call functions $f$ for which there exist finite sets $Q$ and $T$ such that every hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ with $\lambda\in\Lambda_{Q,T}$ computes $f$. There is a naive algorithm (by enumerating all the words for $Q$ and $T$) that permits us to learn the class of guessable functions.
  • For the class of guessable functions, a natural question is whether there is a “fast” algorithm that permits us to learn this class, by making more use of counter-examples, in the style of the successful approaches used for fields and principal ideal domains. These approaches typically alternate between incrementation of the sets $Q$ and $T$, eventually converging on a pair for which $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. A potential problem here is that, in general, one does not know for a given set $Q$ with $\Lambda_{Q}\neq\emptyset$ whether it is worth persevering with incrementation of $T$ or not. We define strongly guessable functions as the ones for which for every suitable set $Q$, there exist a finite set $T$, such that every associated hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ with $\lambda\in\Lambda_{Q,T}$ computes $f$.
@@
In Section 3, we define the classes of weakly guessable, guessable and strongly guessable functions and give several equivalent characterisations for them in Theorem 3.5 and 3.12. We outline a simple learning algorithm for (strongly) guessable functions and discuss further algorithmic issues to be considered including guarantees for termination. In Section 4, we give a general hierarchy for these classes with some specific examples, and give conditions for this hierarchy to collapse. This allows us to have a clearer picture of what happens for several well-studied examples of semirings arising in application areas, where limitations of this approach may be of interest. This opens many questions for further research that are stated in the concluding Section 5.
@@ 3.1. Weakly guessable functions
Observe that by Theorem 2.9, this implies that for all $\lambda\in\Lambda_{Q}$ the automaton $\mathcal{H}_{Q,\lambda}$ computes $f$ and hence for any non-empty $T$, there exists a hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ computing $f$ since $\Lambda_{Q}\subseteq\Lambda_{Q,T}$. So this definition means that there is “some hope” that an algorithm à la Angluin can find a correct automaton, but does not really give a handle on what these functions look like. The following definition turns out to provide a syntactic charaterisation of these functions in terms of automata and links to residual automata. A literal automaton is a residual automaton, as defined for example for non-deterministic or nominal automata [20, 28] with extra conditions to deal with the weights. It is very close to the reduced residual automata in the stochastic settings defined in [15], with a slight difference in the way the weights are handled - the reason being that, while non-zero probabilities have an inverse for the multiplication, this is not systematically the case in general semirings. We use the terminology literal automata to avoid any confusion with [15], our choice of terminology reflecting the fact that each state of a literal automaton can be labelled by a word with the property that any run from the initial state labelled by that word terminates in the given state.
@@
This means that in a literal automaton $\mathcal{A}$, each state $q$ can be characterised by a word $\sigma(q)$ that leads uniquely to that state (and with weight $1_{S}$), and hence the set $\{\langle q\rangle\mid q\text{ is a state of }\mathcal{A}\}$ defined in Theorem 2.9 is exactly the set of rows of the Hankel matrix of $\mathcal{A}$ indexed by the words $\sigma(q)$.
@@
Comparing Theorem 3.5 (characterising weakly guessable functions) with Theorem 2.9 (characterising functions computed by a finite weighted automata), notice that point (3) of Theorem 3.5 states that there is a finite subset of the rows of the Hankel matrix that left-generates every row of the Hankel matrix, whilst Theorem 2.9, states that the rows of the Hankel matrix can be generated by finitely many elements of $S^{\Sigma^{*}}$ (which need not themselves be rows of the Hankel matrix; see also Remark 2.10).
@@
(1. implies 2.) Since $f$ is weakly guessable there is a finite set of rows $Q$ of the Hankel matrix such that $\Lambda_{Q}$ is non-empty; by a slight abuse of notation, we shall view $Q$ as a subset of $\Sigma^{*}$, identifying a row of a Hankel matrix with the word labelling that row. Fix $\lambda\in\Lambda_{Q}$ and recall that (by Theorem  2.9) $\mathcal{H}_{Q,\lambda}$ computes $f$. Let $Q^{\prime}$ denote the set of all prefixes of words in $Q$, and define $\mu=(\mu^{\varepsilon}_{q},\mu_{q,a,p})_{p,q\in Q^{\prime},a\in\Sigma}$ as follows:
  • $\mu^{\varepsilon}_{q}=1_{S}$ if and only if $q=\varepsilon$ and $0_{S}$ otherwise,
  • if $q\in Q^{\prime}$, $a\in\Sigma$ and $qa\in Q^{\prime}$, then $\mu_{q,a,qa}=1_{S}$, and $\mu_{q,a,p}=0_{S}$ for all other $p\in Q^{\prime}$;
@@
  • if $q\in Q$, $a\in\Sigma$ and $qa$ is not in $Q^{\prime}$, then $\mu_{q,a,p}=\lambda_{q,a,p}$ for all $p\in Q$ and $\mu_{q,a,p}=0_{S}$ for all $p\in Q^{\prime}\setminus Q$;
@@
  • if $q$ in $Q^{\prime}\setminus Q$, $a$ in $\Sigma$ and $qa$ is not in $Q^{\prime}$, then for each $p\in Q^{\prime}$ we define $\mu_{q,a,p}$ by induction on the length of $q$, as follows:
@@
    If $q=\varepsilon$, then for all words $w$ we have $\langle\varepsilon\rangle_{w}=\bigoplus_{r\in Q}\lambda^{\varepsilon}_{r}\otimes\langle r\rangle_{w}$ and
    $\langle a\rangle_{w}=\langle\varepsilon\rangle_{aw}=\bigoplus_{r\in Q}\lambda^{\varepsilon}_{r}\otimes\langle r\rangle_{aw}=\bigoplus_{r\in Q}\lambda^{\varepsilon}_{r}\otimes\langle ra\rangle_{w}=\bigoplus_{r,p\in Q}\lambda^{\varepsilon}_{r}\otimes\lambda_{r,a,p}\otimes\langle p\rangle_{w}.$
    In this case we set: $\mu_{\varepsilon,a,p}=\bigoplus_{r\in Q}\lambda^{\varepsilon}_{r}\otimes\lambda_{r,a,p}$ for all $p\in Q$ and $\mu_{\varepsilon,a,p}=0_{S}$ for all $p\in Q^{\prime}\setminus Q$. If $q=ra$ and $\mu_{r,a,p}$ has already been defined for all $p\in Q^{\prime}$ in such a way that expresses $\langle ra\rangle$ as a left-linear combination of the rows $Q$ (i.e. such that $\mu(r,a,p)=0_{S}$ if $p$ is not in $Q$), then for all words $w$ and all $b\in\Sigma$ we have
    $\langle qb\rangle_{w}$ $=$ $\langle rab\rangle_{w}=\langle ra\rangle_{bw}=\bigoplus_{p\in Q}\mu_{r,a,p}\otimes\langle p\rangle_{bw}$
    $=$ $\bigoplus_{p\in Q}\mu_{r,a,p}\otimes\langle pb\rangle_{w}=\bigoplus_{p,s\in Q}\mu_{r,a,p}\otimes\lambda_{p,b,s}\otimes\langle s\rangle_{w}.$
@@
    In this case we set $\mu_{ra,b,p}=\bigoplus_{s\in Q}\mu_{r,a,p}\otimes\lambda_{p,b,s}$ for all $p\in Q$ and $\mu_{ra,b,p}=0_{S}$ for all $p\in Q^{\prime}\setminus Q$.
@@
By construction we then have that $\mu$ belongs to $\Lambda_{Q^{\prime}}$ and hence by Lemma 2.9, $\mathcal{H}_{Q^{\prime},\mu}$ computes $f$. The automaton $\mathcal{H}_{Q^{\prime},\mu}$ is also literal by construction, which concludes the proof.
@@
(2. implies 3.) Let $\mathcal{A}$ be a literal automaton computing $f$ and $Q$ its set of states. By Theorem 2.9, the set $R=\{\langle q\rangle\mid q\in Q\}$, where $\langle q\rangle_{w}$ is defined as the weight of $w$ from $q$, is row-generating for $f$. Moreover, by definition of literal automaton, there is a finite set of words $\{w_{q}\mid q\in Q\}$ such that for all $q$, $\langle q\rangle=\langle w_{q}\rangle$.
@@
(3. implies 4.) Let $W\subseteq\Sigma^{*}$ such that $\{\langle w\rangle\mid w\in W\}$ is row-generating and let a sequence of left-semimodules $(X_{i})_{i\in\mathbb{N}}$ as in the definition of the weak ascending chain condition. Then there is $n$ such that for all $m\geq n$, all the rows $\langle w\rangle$ for $w$ in $W$ are in $X_{m}$, and since they are row-generating, $X_{m}=X_{n}$.
@@
(4. implies 1.) For all positive integers $i$, let $X_{i}$ be the left-semimodule generated by the rows indexed by words of length at most $i$. Then the $X_{i}$ satisfy the hypothesis of the weak acending chain condition and hence there is $n$ such that for all $m$, $X_{m}=X_{n}$. Let $Q$ be the set of words of length at most $n$. Necessarily the set of rows indexed by $Q$ is row-generating (and hence left-closed, and left-generates $\langle\varepsilon\rangle$) so there exists some $\lambda\in\Lambda_{Q}$, which concludes the proof. ∎

### v4

@@ 2.2. Learning with membership and equivalence queries
Suppose that $Q\subseteq S^{\Sigma^{*}}$ is a finite subset such that $\Lambda_{Q}\neq\emptyset$ (or equivalently by Remark 2.7, $Q$ is left-closed and left-generates $\langle\varepsilon\rangle$) and let $\underline{\lambda}\in\Lambda_{Q}$. We shall show that $\mathcal{H}_{Q,\underline{\lambda}}$ computes $f$. To begin with, for each $\underline{q}$ in $Q$, let $\mathcal{H}_{\underline{q}}$ be the automaton obtained from $\mathcal{H}_{Q,\underline{\lambda}}$ by imposing the condition that $\underline{q}$ is the unique initial state with weight $1_{S}$, and keeping all other weights the same. We claim that for all $w$ in $\Sigma^{*}$, $\underline{q}_{w}$ is equal to the weight computed by $\mathcal{H}_{\underline{q}}$ on input $w$. The proof is by induction on the length of $w$. If $w=\varepsilon$, then $\underline{q}_{\varepsilon}$ is (by construction) equal to the final weight of $\underline{q}$ in $\mathcal{H}_{Q,\underline{\lambda}}$ which proves the property. For $a$ in $\Sigma$ and $w$ in $\Sigma^{*}$, $\underline{q}_{aw}=(\underline{q}\cdot a)_{w}=\bigoplus_{\underline{p}\in Q}\underline{\lambda}_{\underline{q},a,\underline{p}}\otimes\underline{p}_{w}$ since $\underline{\lambda}\in\Lambda_{Q}$. By induction hypothesis, $\underline{p}_{w}$ is the weight computed by $\mathcal{H}_{\underline{p}}$ on input $w$, and moreover $\underline{\lambda}_{\underline{q},a,\underline{p}}$ is (by construction) the weight of the transition from state $\underline{q}$ to state $\underline{p}$ labelled by $a$. So $\underline{q}_{aw}$ is indeed the weight computed by $\mathcal{H}_{\underline{q}}$ on input $aw$. To conclude the proof, recall that $\langle\varepsilon\rangle$ denotes the row of $\varepsilon$ in the Hankel matrix of $f$, and so by definition of the Hankel matrix of $f$ and the fact that $\underline{\lambda}\in\Lambda_{Q}$ we have $f(w)=\langle\varepsilon\rangle_{w}=\bigoplus_{\underline{q}\in Q}\underline{\lambda}_{q}\otimes\underline{q}_{w}$; by our previous observation (together with the fact that $\underline{\lambda}_{Q}$ is the initial state vector of $\mathcal{H}_{Q,\underline{\lambda}}$), this is then exactly the value computed by $\mathcal{H}_{Q,\lambda}$ on input $w$.
@@
Now suppose that $f:\Sigma^{*}\to S$ is computed by a finite-state weighted automaton $\mathcal{A}$ over $S$ with set of states $Q$. For each $q\in Q$, let ${}_{q}\mathcal{A}$ (resp. $\mathcal{A}_{q}$) denote the automaton obtained from $\mathcal{A}$ by making $q$ the unique initial (resp. final) state with weight $1_{S}$. For each $q\in Q$, define $\underline{q}\in S^{\Sigma^{*}}$ where for all $u\in\Sigma^{*}$, $\underline{q}_{u}$ is equal to the value computed by ${}_{q}\mathcal{A}$ on input $u$. Let $\underline{Q}=\{\underline{q}\mid q\in Q\}$. First $\underline{Q}$ is left-closed since, by definition, for all words $u\in\Sigma^{*}$, all states $q\in Q$ and all $a\in\Sigma$ we have
$(\underline{q}\cdot a)_{u}=\underline{q}_{au}=\bigoplus_{p\in Q}(\nu_{p}\otimes\underline{p}_{u})$
@@
where $\nu_{p}$ is the weight of the transition labelled by $a$ in $\mathcal{A}$ from $q$ to $p$. Let $w$ be a word in $\Sigma^{*}$ and as before let $\langle w\rangle$ denote the infinite row of the Hankel matrix of $f$ indexed by $w$. Let $\mu_{q}(w)$ be the value computed by $\mathcal{A}_{q}$ on input $w$. Then, by definition, $\langle w\rangle=\bigoplus_{q\in Q}\mu_{q}(w)\otimes\underline{q}$, showing that $\underline{Q}$ is row-generating for $f$ and hence in particular $\langle\varepsilon\rangle$ is left-generated by $\underline{Q}$. ∎
@@
In this section we have taken an approach based on “rows”, describing how to construct hypothesis automata utilising finitely many partial rows of the Hankel matrix. Clearly, a dual construction can be made working with the transpose of the Hankel matrix, that is, using partial columns and corresponding notions of right-linear combinations, right semimodules, right-generating sets, right-closed sets, column-generating sets, etc. In the text that follows we shall refer to the analogues of Hankel automata and hypothesis automata constructed by this dual method as co-Hankel automata and co-hypothesis automata. For the readers convenience we outline the details of these constructions and summarise the dual definition to those presented in the main text in Appendix A.
@@
It is known that for all functions computed by weighted automata over fields there is an efficient algorithm (using linear algebra) to find, by membership and equivalence queries, finite sets of words $Q$ and $T$ such that $\Lambda_{Q,T}$ is non-empty together with an element $\underline{\lambda}\in\Lambda_{Q,T}$ such that the hypothesis automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ computes the target function (see for example [9]). For an arbitrary (but fixed) semiring $S$, the following questions arise:
@@
  • For which functions $f:\Sigma^{*}\rightarrow S$ does there exist an hypothesis automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ computing $f$? We will call the class of functions which can be computed by at least one hypothesis automaton (which recall is built using a finite portion of the Hankel matrix) weakly guessable. Clearly, if $f$ is not weakly guessable then any learning algorithm based around making hypotheses of the kind under discussion (constructed from certain left linear combinations of partial rows of the Hankel matrix) will fail to learn $f$.
@@
  • For weakly guessable functions over $S$, is there a learning algorithm using membership and equivalence queries that allows a learner to construct a finite sequence of hypothesis automata, where the last in the sequence computes the target function? In other words: when there exists at least one correct hypothesis, can we devise a strategy to find one? A possible issue here is that even when we have settled upon a suitable set $Q$ with $\Lambda_{Q}\neq\emptyset$, it may be that there is no finite set $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}$, and therefore no obvious general method to discover a “correct” $\underline{\lambda}\in\Lambda_{Q}$ in finite time. We say that $f$ is guessable if there exist finite sets $Q$ and $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$ and so every hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ with $\lambda\in\Lambda_{Q,T}$ computes $f$. Since $\Sigma^{*}$ is countable, so too is the set of all pairs of finite subsets of $\Sigma^{*}$; thus if armed with an enumeration of all such pairs together with the ability to compute an element of every left solution set $\Lambda_{Q,T}$ (or else determine that the set is empty), it is clear that there is a naive algorithm (working through the list of all pairs of subsets, and performing an equivalence query where possible) that is guaranteed to succeed on the class of guessable functions (since $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$ holds for at least one pair).
  • In order to find a correct hypothesis automaton, some of the existing approaches and algorithms typically alternate between ‘incrementation’ of the sets of words $Q$ and $T$ adding elements to $Q$ in a prescribed manner when the left-solution set $\Lambda_{Q,T}$ is found to be empty, and adding elements to $T$ in a prescribed manner when a counter example is given, eventually converging on a pair for which $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. A potential problem here is that, in general, one does not know for a given set $Q^{\prime}$ whether it is worth persevering by continuing to add words to $T$ or not (in other words, even though there exists a finite set $Q$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$ for some finite set $T$, it may be the case that for the current choice $Q^{\prime}$ we have $\Lambda_{Q^{\prime},T}\supsetneq\Lambda_{Q^{\prime}}$ for all finite sets $T$). With this in mind, we say that a guessable function $f$ is strongly guessable if for all finite sets of words $Q$ there exists a finite set of words $T$, such that $\Lambda_{Q,T}=\Lambda_{Q}$. For strongly guessable functions, the general problem therefore reduces to the question of how to find a set $Q$ with the property that $\Lambda_{Q}\neq\emptyset$.
  • Finally, given a partial Hankel matrix one could ask the corresponding questions concerning existence of co-hypothesis automata constructed from right-linear combinations of partial columns of the Hankel matrix (see Appendix A for details), leading to the dual notion of weakly co-guessable, co-guessable, and strongly co-guessable functions.
@@
In Section 3, we define the classes of weakly guessable, guessable and strongly guessable functions and give several equivalent characterisations for them in Theorem 3.6 and 3.13. In Section 4, we give a general hierarchy for these classes with some specific examples, and give conditions for this hierarchy to collapse. This allows us to have a clearer picture of what happens for several well-studied examples of semirings arising in application areas, where limitations of this approach may be of interest. This opens many questions for further research that are stated in the concluding Section 5.
@@ 3.1. Weakly guessable functions
Observe that by Theorem 2.12, this implies that for all $\underline{\lambda}\in\Lambda_{Q}$ the Hankel automaton $\mathcal{H}_{Q,\underline{\lambda}}$ computes $f$ and hence for any non-empty set of words $T$, there exists an hypothesis automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ computing $f$ since (by Remark 2.8) $\Lambda_{Q}\subseteq\Lambda_{Q,T}$. So this definition means that there is “some hope” that an algorithm à la Angluin can find a correct automaton, but does not really give a handle on what these functions look like. The following definition (literal automaton) turns out to provide a syntactic charaterisation of these functions in terms of automata, and links to residual automata. A literal automaton is a residual automaton, as defined for example for non-deterministic or nominal automata [22, 27] with extra conditions to deal with the weights. It is very close to the reduced residual automata in the stochastic settings defined in [17], with a slight difference in the way the weights are handled - the reason being that, while non-zero probabilities have an inverse for the multiplication, this is not systematically the case in general semirings. We use the terminology literal automata to avoid any confusion with [17], our choice of terminology reflecting the fact that each state of a literal automaton can be labelled by a word with the property that any run from the initial state labelled by that word terminates in the given state.
@@
Remark 3.3.
In a literal automaton $\mathcal{A}$ with set of states $Q$, each state $q$ can be labelled by a word $\sigma(q)$ which when taken as input terminates in state $q$ (via a unique run, which has weight $1_{S}$), and so in this case the set $\underline{Q}=\{\underline{q}\mid q\text{ is a state of }\mathcal{A}\}$ defined in Theorem 2.12 is easily seen to be precisely the set of rows of the Hankel matrix of $\mathcal{A}$ indexed by the words chosen to label the states, that is, $\underline{Q}=\{\langle\sigma(q)\rangle:q\in Q\}$.
@@
Before presenting the proof, it is instructive to compare Theorem 3.6 (characterising weakly guessable functions) with Theorem 2.12 (characterising functions computed by a finite weighted automata). Notice that point (3) of Theorem 3.6 states that there is a finite subset of the rows of the Hankel matrix that left-generates every row of the Hankel matrix, whilst Theorem 2.12, states that the rows of the Hankel matrix can be generated by finitely many elements of $S^{\Sigma^{*}}$ (which need not themselves be rows of the Hankel matrix; see also Remark 2.13).
@@
(1. implies 2.) Since $f$ is weakly guessable there is a finite set of rows $Q$ of the Hankel matrix such that $\Lambda_{Q}$ is non-empty. Let $W$ be a finite set of words with $|W|=|Q|$ such that $Q$ is the set of rows of the Hankel matrix indexed by $W$ and let $W^{\prime}$ denote the set of all prefixes of words in $W$, with $Q^{\prime}$ the corresponding set of rows of the Hankel matrix, so that $Q\subseteq Q^{\prime}=\{\langle q\rangle\mid q\in W^{\prime}\}$. Since $\Lambda_{Q}\neq\emptyset$, we have seen (in Lemma 2.9) that $\Lambda_{Q^{\prime}}\neq\emptyset$; we aim to show that there exists $\underline{\mu}\in\Lambda_{Q^{\prime}}$ such that the corresponding Hankel automaton $\mathcal{H}_{Q^{\prime},\underline{\mu}}$ (which by Theorem 2.12 computes $f$) is literal. To this end, we let $\underline{\lambda}\in\Lambda_{Q}$ so that (by Theorem  2.12) $\mathcal{H}_{Q,\underline{\lambda}}$ computes $f$ and use the values in $\underline{\lambda}$ to define $\underline{\mu}\in S^{Q^{\prime}\cup(Q^{\prime}\times\Sigma\times Q^{\prime})}$ as follows:
  • $\underline{\mu}_{\langle q\rangle}=1_{S}$ if and only if $\langle q\rangle=\langle\varepsilon\rangle$ and $\underline{\mu}_{\langle q\rangle}=0_{S}$ for all other $\langle q\rangle\in Q^{\prime}$,
  • if $q\in W^{\prime}$, $a\in\Sigma$ and $qa\in W^{\prime}$, then $\underline{\mu}_{\langle q\rangle,a,\langle qa\rangle}=1_{S}$, and $\underline{\mu}_{\langle q\rangle,a,\langle p\rangle}=0_{S}$ for all other $\langle p\rangle\in Q^{\prime}$;
@@
  • if $q\in W$, $a\in\Sigma$ and $qa$ is not in $W^{\prime}$, then $\underline{\mu}_{\langle q\rangle,a,\langle p\rangle}=\underline{\lambda}_{\langle q\rangle,a,\langle p\rangle}$ for all $p\in W$ and $\underline{\mu}_{\langle q\rangle,a,\langle p\rangle}=0_{S}$ for all $\langle p\rangle\in Q^{\prime}\setminus Q$;
@@
  • if $q\in W^{\prime}\setminus W$, $a$ in $\Sigma$ and $qa$ is not in $W^{\prime}$, then for each $\langle p\rangle\in Q^{\prime}$ we define $\underline{\mu}_{\langle q\rangle,a,\langle p\rangle}$ by induction on the length of $q$, as follows:
@@
    If $q=\varepsilon$, then since $\underline{\lambda}\in\Lambda_{Q}$ for all words $w$ we have $\langle\varepsilon\rangle_{w}=\bigoplus_{\langle r\rangle\in Q}\underline{\lambda}_{\langle r\rangle}\otimes\langle r\rangle_{w}$ and so
    $\langle a\rangle_{w}=\langle\varepsilon\rangle_{aw}=\bigoplus_{\langle r\rangle\in Q}\underline{\lambda}_{\langle r\rangle}\otimes\langle r\rangle_{aw}=\bigoplus_{\langle r\rangle\in Q}\underline{\lambda}_{\langle r\rangle}\otimes\langle ra\rangle_{w}=\bigoplus_{\langle r\rangle,\langle p\rangle\in Q}\underline{\lambda}_{\langle r\rangle}\otimes\underline{\lambda}_{\langle r\rangle,a,\langle p\rangle}\otimes\langle p\rangle_{w},$
@@
    using the fact that $\langle ra\rangle=\langle r\rangle\cdot a$. In this case we set: $\underline{\mu}_{\langle\varepsilon\rangle,a,\langle p\rangle}=\bigoplus_{\langle r\rangle\in Q}\underline{\lambda}_{\langle r\rangle}\otimes\underline{\lambda}_{\langle{r}\rangle,a,\langle p\rangle}$ for all $\langle p\rangle\in Q$ and $\mu_{\langle\varepsilon\rangle,a,\langle p\rangle}=0_{S}$ for all $\langle p\rangle\in Q^{\prime}\setminus Q$.
    If $q=ra$ and $\underline{\mu}_{\langle r\rangle,a,\langle p\rangle}$ has already been defined for all $\langle p\rangle\in Q^{\prime}$ in such a way that expresses $\langle q\rangle=\langle ra\rangle=\langle r\rangle\cdot a$ as a left-linear combination of the rows $Q$ (i.e. such that $\underline{\mu}_{\langle r\rangle,a,\langle p\rangle}=0_{S}$ if $\langle p\rangle$ is not in $Q$), then for all words $w$ and all $b\in\Sigma$ we have
    $(\langle q\rangle\cdot b)_{w}=\langle qb\rangle_{w}$ $=$ $\langle rab\rangle_{w}=\langle ra\rangle_{bw}=\bigoplus_{\langle p\rangle\in Q}\underline{\mu}_{\langle r\rangle,a,\langle p\rangle}\otimes\langle p\rangle_{bw}$
    $=$ $\bigoplus_{\langle p\rangle\in Q}\underline{\mu}_{\langle r\rangle,a,\langle p\rangle}\otimes\langle pb\rangle_{w}=\bigoplus_{\langle p\rangle,\langle s\rangle\in Q}\underline{\mu}_{\langle r\rangle,a,\langle p\rangle}\otimes\underline{\lambda}_{\langle p\rangle,b,\langle s\rangle}\otimes\langle s\rangle_{w}.$
    In this case we set $\underline{\mu}_{\langle ra\rangle,b,\langle s\rangle}=\bigoplus_{\langle p\rangle\in Q}\underline{\mu}_{\langle r\rangle,a,\langle p\rangle}\otimes\underline{\lambda}_{\langle p\rangle,b,\langle s\rangle}$ for all $\langle s\rangle\in Q$ and $\mu_{\langle ra\rangle,b,\langle s\rangle}=0_{S}$ for all $\langle s\rangle\in Q^{\prime}\setminus Q$.
@@
By construction we then have that $\underline{\mu}$ belongs to $\Lambda_{Q^{\prime}}$ and hence by Lemma 2.12, $\mathcal{H}_{Q^{\prime},\underline{\mu}}$ computes $f$. The automaton $\mathcal{H}_{Q^{\prime},\underline{\mu}}$ is also literal by construction, which concludes the proof.
@@
(2. implies 3.) Let $\mathcal{A}$ be a literal automaton computing $f$ and $Q$ its set of states. By Theorem 2.12, the set $\underline{Q}=\{\underline{q}\mid q\in Q\}$, where $\underline{q}_{w}$ is defined as the value computed by the automaton $\mathcal{A}_{q}$ on input $w$, is row-generating for $f$ and, as observed in Remark 3.3, there is a finite set of words $\{\sigma(q)\mid q\in Q\}$ such that $\underline{Q}=\{\langle\sigma(q)\rangle\mid q\in Q\}$.
@@
(3. implies 4.) Let $W\subseteq\Sigma^{*}$ be a finite set such that $\{\langle w\rangle\mid w\in W\}$ is row-generating and consider a sequence of left-semimodules $(X_{i})_{i\in\mathbb{N}}$ as in the definition of the weak ascending chain condition. Then there exists $n$ such that for all $m\geq n$, all the rows $\langle w\rangle$ for $w$ in $W$ are in $X_{m}$, and since these elements are row-generating for $f$ it follows that $X_{m}=X_{n}$.
@@
(4. implies 1.) For all positive integers $i$, let $X_{i}$ be the left-semimodule generated by the set $Q_{i}$ containing all rows of $F$ indexed by words of length at most $i$. Then the $X_{i}$ satisfy the hypothesis of the weak acending chain condition and hence there is $n$ such that for all $m$, $X_{m}=X_{n}$. Necessarily the set $Q_{n}$ is row-generating and hence (by Lemma 2.4) left-closed, and left-generates $\langle\varepsilon\rangle$, so there exists some $\lambda\in\Lambda_{Q}$, giving $\Lambda_{Q_{n}}\neq\emptyset$ wich demonstrates that $f$ is weakly guessable and concludes the proof. ∎